# End-to-End Reinforcement Learning of Koopman Models for eNMPC of an Air Separation Unit

### v1

@@ 1 Introduction
Data-driven dynamic models can be trained in an end-to-end fashion for optimal performance as part of (economic) (nonlinear) model predictive control ((e)(N)MPC) (e.g., gros2019data; amos2018differentiable). We recently introduced a method (mayfrank2024end) based on reinforcement learning (RL) for end-to-end learning of Koopman surrogate models (koopman1931hamiltonian; korda2018linear) for (e)NMPC applications. Such data-driven surrogate models can make eNMPC computationally tractable in case an accurate mechanistic process model is available but using it as part of a model predictive controller is too computationally expensive. Moreover, in scenarios where no reliable model is available, the same framework can learn directly from plant data. Alternative methods for end-to-end learning of data-driven models for control applications focus on linear models (chen2019gnu), optimize highly-structured models with few parameters requiring expert system knowledge (brandner2023reinforcement), cannot handle hard constraints on system states (amos2018differentiable), or are only applicable to setpoint tracking problems (iwata2022data; yin2022embedding). Our method can optimize highly parameterized and thus flexible Koopman models for control problems with state constraints and arbitrary convex objective functions.
@@ 4 Numerical experiments
Fig. 4 illustrates the evolution of the reward over 200,000 environment steps for each training run. Each training run takes approx. two days on a Windows 11 workstation with an Intel Core i7-14700 CPU. After training has completed, our final performance evaluation is based on the policy that attained the highest average reward during a policy rollout. Consequently, our primary metric of interest is the maximum average reward achieved by each training run, rather than the reward at the final training step. Three out of five training runs demonstrate a substantial improvement in the policy’s average reward compared to the initial guess.
@@ 5 Conclusion
Among the five independent end-to-end training runs, three significantly improve eNMPC performance over the SI baseline, reducing and even eliminating constraint violations while maintaining high economic efficiency.

### v2

@@ 1 Introduction
Data-driven dynamic models can be trained in an end-to-end fashion for optimal performance as part of (economic) (nonlinear) model predictive control ((e)(N)MPC) (e.g., gros2019data; amos2018differentiable). We recently introduced a method (mayfrank2024end) based on reinforcement learning (RL) for end-to-end learning of Koopman surrogate models (korda2018linear) for (e)NMPC applications. Such data-driven surrogate models can make eNMPC computationally tractable in case an accurate mechanistic process model is available but using it as part of a model predictive controller is too computationally expensive. Moreover, in scenarios where no reliable model is available, the same framework can learn directly from plant data. Alternative methods for end-to-end learning of data-driven models for control applications focus on linear models (chen2019gnu), optimize highly-structured models with few parameters requiring expert system knowledge (brandner2023reinforcement), cannot handle hard constraints on system states (amos2018differentiable), or are only applicable to setpoint tracking problems (iwata2022data). Our method can optimize highly parameterized and thus flexible Koopman models for control problems with state constraints and arbitrary convex objective functions.
@@ 4 Numerical experiments
Fig. 4 illustrates the evolution of the reward over 200,000 environment steps for each training run. Each training run takes approx. two days on a Windows 11 workstation with an Intel Core i7-14700 CPU. Three elements contribute to overall training time: (i) simulation of the mechanistic DAE environment, (ii) repeated solution of the convex eNMPC problem, and (iii) gradient-based policy optimization. The dominant contributor to the training time is the policy optimization step, which requires backpropagating through the Koopman eNMPC. In contrast, pure policy inference typically takes well under one second per control step, making the controller easily real-time capable. After training has completed, our final performance evaluation is based on the policy that attained the highest average reward during a policy rollout. Consequently, our primary metric of interest is the maximum average reward achieved by each training run, rather than the reward at the final training step. Three out of five training runs demonstrate a substantial improvement in the policy’s average reward compared to the initial guess.
@@ 5 Conclusion
Among the five independent end-to-end training runs, three significantly improve eNMPC performance over the SI baseline, reducing and even eliminating constraint violations while maintaining high economic efficiency. Compared to our earlier two-state CSTR case study (mayfrank2024end), where all ten training runs consistently improved performance and rewards saturated after roughly 180,000 steps, the present large-scale ASU problem poses a more challenging learning task, reflected in the lower fraction of successful runs. Note that while our method strongly incentivizes constraint satisfaction, it does not provide formal theoretical guarantees.